Dresden Pharmaceuticals has decided to go ahead and start clinical trials on a potential new drug. The total R&D costs are estimated to reach around $875,000,000 with clinical trials mounting to $145,000,000. The current market size is estimated to be around 3,000,000 and is expected to grow at 4 percent every year. The market share Dresden hopes to capture in the first year is 7 percent, and is projected to grow by 25 percent each year for the next 4 years. A monthly prescription is anticipated to generate revenue of $420 while incurring variable costs of $150. A discount rate of 8 percent is assumed.

Dresden Pharmaceuticals
Data
Market Size 3,000,000
Unit (monthly Rx) revenue ($) 420
Unit (monthly Rx) cost ($) 150
Discount Rate ( per cent) 8
Project Costs
R&D ($) 875,000,000
Clinical Trials ($) 145,000,000



Calculate the annual cost incurred for the second year.

A. $638,820,000
B. $1,224,120,000
C. $884,520,000

D. $491,400,000

Which of the following conditions is the optimal solution to the newsvendor problem, where Q is the quantity to be purchased, and D is demand?

A. Q / D = 0
B. Q > D
C. Q = D
D. D > Q

Calculate cumulative net profit at the fourth year.

A. $1,073,538,462
B. $3,189,634,800
C. $1,494,838,800
D. $1,312,041,240
Answer #1

market size for first year=3000000

market size in second year =3000000*1.04=3120000

market captured in first year =7%

market captured in second year=7%*1.25=8.75%

no of units in second year=8.75%*312000=273000

cost of total units=273000*150*12=491400000(d)

2.q>d becausethere iss always a scope for higher demand for the product than the meanso hence if we stock exactly the same amount as the mean demand then we might not be able to supply for the extra demand.

3.market size for fourth year =3000000*(1.04)^4=3509575

market cap for fourth year=7*(1.25)^4=17.089%

no of units in fourth year=17.089*3509575=59978088

net profit=59978088*(420-150)*12=1943290063